How do I solve the following:
When Sally asked Johnny how old he was, he replied that in 2 years he would be twice as old as he was 3 years ago. How old is Johnny?

Answer by rothauserc(4718)

let x be Johnny's are
x + 2 = 2(x-3)
x+2 = 2x - 6
x = 8
Johnny is 8